OpenStudy (anonymous):

Write the equation of the circle with center (-2, -5) and containing the point (-10, -20). Use the ^ key for the exponents. can some one help?

OpenStudy (anonymous):

first find the radius by using the distance formula. find the distance between (-2, -5) and (-10, -20)... what u got?

OpenStudy (anonymous):

(-8, -15)?

OpenStudy (anonymous):

no.... distance....

OpenStudy (anonymous):

Distance = \(\large \sqrt{(x_2-x_1)^2+(y_2-y_1)^2} \)

OpenStudy (anonymous):

289?

OpenStudy (anonymous):

actually you'd square it so 17 right?

OpenStudy (anonymous):

yes... so your radius is 17...

OpenStudy (anonymous):

now the equation of a circle with center (h, k) is written in standard form as: \(\huge (x-h)^2+(y-k)^2=r^2 \)

OpenStudy (anonymous):

can you plug in your center and radius and then simplify... what u got?

OpenStudy (anonymous):

how do you get the center?

OpenStudy (anonymous):

the center is given....

OpenStudy (anonymous):

center (-2, -5) radius = 17 put in standard form....

OpenStudy (anonymous):

(-10-2)^2 +(-20-5)^2=17^2

OpenStudy (anonymous):

so it would be (-12)^2+ (-25)^2=17^2